# Revisiting Shooting Point Monte Carlo Methods for Transition Path Sampling

### v2

@@ II.7 Flexible-length Aimless Shooting
Figure 3: The two-point aimless shooting move in the extended space of paths $X$ and shooting indices $k$. (1) The move starts with the current path $X^{\text{o}}$, shooting index $k^{\text{o}}$ and direction $s^{\text{o}}$ that determines the position of the valid shooting indices. (2) The first index shift is performed by selecting one of the two potential shooting indices which generates $\hat{k}^{\text{o}}$. (3) An extended space shooting move generates $X^{\text{n}}$ and $\hat{k}^{\text{n}}$. At the same time, a new direction $s^{\text{n}}$ is drawn that determines if the second potential shooting point is on the forward or backward segment of the new path. (4) A second index shift is performed leading to $k^{\text{n}}$. This shift is required to ensure the reverse move is possible (tracing bottom to top).
@@
As mentioned above, there is a crucial difference between the fixed-length and the flexible-length moves. Namely, in flexible-length algorithm, the new shooting index $k^{\text{n}}$ is selected depending on the previous index $k^{\text{o}}$ (see Fig. 3). In particular, the two-point variant described in Mullen et al. (2015) samples the new shooting index from $\{k^{\text{o}},k^{\text{o}}+s\Delta k\}$, where $s\in\{-1,1\}$ defines whether the second shooting point is on the backward or the forward trajectory segment. Since the sign $s$ is preserved until a next trial is accepted, we include it in the extended ensemble $(X,k,s)$ (see SI for detailed balance derivation). To respect reversibility, the generation probability of $k^{\text{n}}$ is divided into two steps. We first choose a shooting index $\hat{k}^{\text{o}}$ on the old path according to
@@
The new path $X^{\text{n}}$ is generated by shooting from $\hat{k}^{\text{o}}$ and integrating the equations of motion until a stable state is reached. We then obtain a new sign $s^{\text{n}}$ by drawing from $\{-1,1\}$ with equal probability. We assume $p(s^{\text{n}})=p_{\text{gen}}(s^{\text{n}})=\frac{1}{2}$. Denoting $\hat{k}^{\text{n}}$ as the index of the shooting point on $X^{\text{n}}$, the new shooting index $k^{\text{n}}$ is generated according to:
$p_{\text{gen}}(k^{\text{n}}|$ $X^{\text{n}},\hat{k}^{\text{n}},s^{\text{n}},X^{\text{o}},k^{\text{o}})=$
$\begin{cases}\frac{1}{2}\ \text{if }k^{\text{n}}\in\{\hat{k}^{\text{n}},\hat{k}^{\text{n}}+s^{\text{n}}\Delta k\},\\ 0\ \text{else}\end{cases}\,.$ (28)
This two-step procedure ensures that the path $X^{\text{o}}$ and shooting index $k^{\text{o}}$ can be generated from $\{X^{\text{n}}$, $k^{\text{n}}\}$ in the reverse move. To simplify the notation, we summarize the $1/2$ generation probabilities from Eqs. (27), (II.7) and $p_{\text{gen}}(s)$ as:
$p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}},s^{\text{n}})\,p_{\text{gen}}(s^{\text{n}})$
$=p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}},s^{\text{o}})\,p_{\text{gen}}(s^{\text{o}})=\frac{1}{8}\,.$ (29)
When inserting these generation probabilities into the analogous form of Eq. II.3 derived in the SI, we obtain
@@
Hence, we are left with the ratio of the shooting index distributions, which are not included in the original algorithm Mullen et al. (2015). As a result, accepting a path if $H_{\text{AB}}(X^{\text{n}})=1$ in a flexible-length setting violates detailed balance, which is also apparent in the results of numerical simulations (see section III). However, following the described two-step procedure, a properly weighted path ensemble can be obtained even a posteriori by assigning each path a statistical weight of $\Omega(X)=1/L(X)$ under the assumption of a uniform shooting index distribution (see SI for derivation).
@@ III Results and Discussion
Within the flexible-length path ensemble, we observe that uniform two-way shooting converges to the reference ensemble, as indicated by overlapping $\rho(x|\text{TP})$ distributions and path length distributions $p(L)$ (Fig. 6). The density of points at $k$ denoted as $\rho_{\text{ens}}(x)$ matches the reweighted density of points on paths (Eq. 15) obtained from the reference trajectory since shooting indices are selected with a uniform $p(k|X)$ on each path. In contrast, flexible-length aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) do not converge to the reference, as shown in panels A and C of Fig. 6. Panel B shows that the shooting points for spring shooting are much more focused around the location of the barrier. While this is a desired feature, it clearly gives rise to the wrong distribution of transition paths (panels A and C). We note that the deviations from the reference path ensemble are less apparent at higher barriers, where the distribution of shooting points is more localized for all methods. This is also the reason why spring shooting appeared to produce seemingly converged results in Brotzakis and Bolhuis (2016).
Including the shooting index distribution $p(k|X)$ in the acceptance probability and, for spring shooting, performing a second shift of the shooting index, recovers the convergence to the reference path ensemble (panels D, E, and F of Fig. 6). However, since we impose a uniform $p(k|X)$, shooting moves are now initiated from the same points as in standard uniform two-way shooting. In addition, the small displacements of the shooting index in aimless and spring shooting introduce correlations between shooting points that are absent in uniform one- or two-way shooting. These correlations result in slow exploration of shooting points away from the barrier top. For sufficiently large barriers, this can cause disproportionate shooting from the barrier top during the early stages of sampling, which in turn leads to a slow convergence to the correct path ensemble.
@@ Acknowledgements.
This research was funded in part by the Austrian Science Fund (FWF) 10.55776/F81. For Open Access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
@@
Sebastian Falkner1,2, Alessandro Coretti1, Baron Peters3,4, Peter G. Bolhuis5, and Christoph Dellago1,∗
1Faculty of Physics, University of Vienna, 1090 Vienna, Austria.
@@
Since this should be true for all values of k, we require that the marginal of $P(X,k)$ in the reweighted flexible aimless path ensemble will be equal to the marginal of $P(X,k)$ in the correct ensemble That is,

### v3

@@ II.7 Flexible-length Aimless Shooting
Figure 3: The two-point aimless shooting move in the extended space of paths $X$ and shooting indices $k$. (1) The move starts with the current path $X^{\text{o}}$, shooting index $k^{\text{o}}$ and direction $s^{\text{o}}$ that determines the position of the valid shooting indices. (2) The first index shift is performed by selecting one of the two potential shooting indices which generates $\hat{k}^{\text{o}}$. The direction $\hat{s}^{\text{o}}$ is flipped as the alternative shooting index is selected. (3) An extended space shooting move generates $X^{\text{n}}$ and $\hat{k}^{\text{n}}$. At the same time, a new direction $\hat{s}^{\text{n}}$ is drawn that determines if the second potential shooting point is on the forward or backward segment of the new path. (4) A second index shift is performed leading to $k^{\text{n}}$ with corresponding direction $s^{\text{n}}$. This shift is required to ensure the reverse move is possible (tracing bottom to top).
@@
As mentioned above, there is a crucial difference between the fixed-length and the flexible-length moves. Namely, in flexible-length algorithm, the new shooting index $k^{\text{n}}$ is selected depending on the previous index $k^{\text{o}}$ (see Fig. 3). In particular, the two-point variant described in Mullen et al. (2015) samples the new shooting index from $\{k^{\text{o}},k^{\text{o}}+s^{\text{o}}\Delta k\}$, where $s^{\text{o}}\in\{-1,1\}$ defines whether the second shooting point is on the backward or the forward trajectory segment. Since the sign $s$ is preserved until a next trial is accepted, we include it in the extended ensemble $(X,k,s)$ (see SI for detailed balance derivation). To respect reversibility, the generation probability of $k^{\text{n}}$ is divided into two steps. We first choose a shooting index $\hat{k}^{\text{o}}$ on the old path according to
@@
If the opposing point, namely $k^{\text{o}}+s^{\text{o}}\Delta k$ is selected, we flip the direction $s^{\text{o}}$ such that the two possible shooting indices in the new state are preserved (see Fig 3 at step 2). The new path $X^{\text{n}}$ is generated by shooting from $\hat{k}^{\text{o}}$ and integrating the equations of motion until a stable state is reached. We then obtain a new sign $\hat{s}^{\text{n}}$ by drawing from $\{-1,1\}$ with equal probability. We assume $p(\hat{s}^{\text{n}})=p_{\text{gen}}(\hat{s}^{\text{n}})=\frac{1}{2}$. Denoting $\hat{k}^{\text{n}}$ as the index of the shooting point on $X^{\text{n}}$, the new shooting index $k^{\text{n}}$ is generated according to:
$p_{\text{gen}}(k^{\text{n}}|$ $X^{\text{n}},\hat{k}^{\text{n}},\hat{s}^{\text{n}})=\begin{cases}\frac{1}{2}\ \text{if }k^{\text{n}}\in\{\hat{k}^{\text{n}},\hat{k}^{\text{n}}+\hat{s}^{\text{n}}\Delta k\},\\ 0\ \text{else}\end{cases}\,.$ (28)
Again, we obtain $s^{\text{n}}$ by flipping $\hat{s}^{\text{n}}$ if the opposing point was selected or else we leave it unmodified. In case $\hat{k}^{\text{o}}$ or $k^{\text{n}}$ fall outside of the old or new path, respectively, the move is rejected and the old path and shooting index are counted again.
This two-step procedure ensures that the path $X^{\text{o}}$ and shooting index $k^{\text{o}}$ can be generated from $\{X^{\text{n}}$, $k^{\text{n}}\}$ in the reverse move. We note that generation probabilities in Eqs. (27), (28) and $p_{\text{gen}}(s)$ are symmetric with probability $1/2$ and, as a result, they cancel in the acceptance criterion (Eq. II.3). In that case, given the analogous form of Eq. II.3 derived in the SI, we obtain
@@
Hence, we are left with the ratio of the shooting index distributions, which are not included in the original algorithm Mullen et al. (2015). As a result, accepting a path if $H_{\text{AB}}(X^{\text{n}})=1$ in a flexible-length setting violates detailed balance, which is also apparent in the results of numerical simulations (see section III). In the following, we choose a uniform shooting index distribution $p(k|X)=1/L(X)$ as it requires no prior knowledge of the simulated system such as an order parameter. In that case, following the described two-step procedure, a properly weighted path ensemble can be obtained even a posteriori by assigning each path a statistical weight of $\Omega(X)=1/L(X)$ in the case of a uniform shooting index distribution (see SI for derivation).
@@ III Results and Discussion
Within the flexible-length path ensemble, we observe that uniform two-way shooting converges to the reference ensemble, as indicated by overlapping $\rho(x|\text{TP})$ distributions and path length distributions $p(L)$ (Fig. 6). The density of points at $k$ denoted as $\rho_{\text{ens}}(x)$ matches the reweighted density of points on paths (Eq. 17) obtained from the reference trajectory since shooting indices are selected with a uniform $p(k|X)$ on each path. In contrast, flexible-length aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) do not converge to the reference, as shown in panels A and C of Fig. 6. Panel B shows that the shooting points for spring shooting are much more focused around the location of the barrier. While this is a desired feature, it clearly gives rise to the wrong distribution of transition paths (panels A and C). We note that the deviations from the reference path ensemble are less apparent at higher barriers (see SI), where the distribution of shooting points is more localized for all methods. This is also the reason why spring shooting appeared to produce seemingly converged results in Brotzakis and Bolhuis (2016).
Including the shooting index distribution $p(k|X)$ in the acceptance probability and, for spring shooting, performing a second shift of the shooting index, recovers the convergence to the reference path ensemble (panels D, E, and F of Fig. 6). However, even though the methods differ in the way a new shooting index is drawn in each trial, shooting moves are now initiated from the same points as in standard uniform two-way shooting since we impose the same shooting index distribution. In addition, the small displacements of the shooting index in aimless and spring shooting introduce correlations between shooting points that are absent in uniform one- or two-way shooting. These correlations result in slow exploration of shooting points away from the barrier top (see SI). For flexible length aimless shooting and spring shooting, the shooting points were also generated disproportionately near the barrier regions. To maintain this desired focus, small displacements of the shooting points were performed in successive trials. This behavior is partially retained in the revised algorithms in the early stages of sampling, as the initial shooting index is often selected near the barrier top. However, with the imposed uniform shooting index distribution, the shooting index occasionally drifts away from the barrier to ensure full sampling the entire index distribution. This leads to low acceptance probabilities and extended sequences of rejected pathways, until the index returns to the barrier top by multiple small displacements in the right direction. The resulting persistent correlations considerably lower the efficiency of the algorithm. In standard TPS, on the other hand, the shooting index is redrawn independently at each trial, allowing it to move away from and return to the barrier instantaneously without such correlations.
@@ Acknowledgements.
This research was funded in part by the Austrian Science Fund (FWF) 10.55776/F8100. For Open Access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
@@
Sebastian Falkner1,2, Alessandro Coretti1, Baron Peters3,4, Peter G. Bolhuis5, and Christoph Dellago1,6,∗
1University of Vienna, Faculty of Physics, 1090 Vienna, Austria.
@@
6 Research Platform on Accelerating Photoreaction Discovery (ViRAPID), University of Vienna, 1090 Vienna, Austria.
@@
Since this should be true for all values of $k$, we require that the marginal of $P_{\mathrm{AB}}(X,k)$ in the reweighted flexible aimless path ensemble will be equal to the marginal of $P_{\mathrm{AB}}(X,k)$ in the correct ensemble That is,
@@
Convergence of Path Sampling Algorithms
Figure S1: (A) Convergence of flexible-length transition path sampling algorithms including uniform two-way shooting, aimless shooting as proposed by Mullen et al. Mullen et al. (2015), spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) and both with the proposed corrections. We use the Wasserstein distance between the reference distribution of points on paths $p_{\mathrm{ref}}(x|\mathrm{TP})$ and the distribution from each method at a specific trial $p_{\mathrm{TP}}(x|\mathrm{TP})$ averaged over $48$ runs. (B) Wasserstein distance between the distributions obtained from long simulations using the unrevised spring and aimless algorithms and the reference distribution as a function of the barrier height. Each point represents the average distance after $10^{6}$ trials from $192$ simulations. (C) Wasserstein distance with respect to the reference path length distribution for different barrier heights.
In the main text, we show that the original spring and aimless algorithms do not converge to the correct distribution of points on paths. Here, we provide more details on the rate of convergence and the error of the unrevised methods with respect to the barrier height for the toy system (Fig. S1). We measure convergence numerically using the Wasserstein distance between the reference distribution of points on the paths and the distribution generated by each method after a certain number of trials (Fig. S1A). The convergence rate of the corrected methods is slower compared to regular two-way shooting with a uniform shooting index selection. This is, as stated in the main text, due to the introduction of correlations in the selection process of the new shooting point. Once the old shooting point $k^{\mathrm{o}}$ fluctuates away from the top of the barrier, the acceptance rate decreases drastically and the convergence is slowed. We see that the error of the original algorithms varies with the barrier height of the system (Fig. S1B), where for spring shooting we see an initial increase in the error followed by a continuous decrease. For the original aimless algorithm, the error decreases initially with barrier height, however, at high barriers the error increases and the low acceptance rate leads to slow convergence. The error in the path length distribution continuously decreases with increasing barrier heights (Fig. S1C), up to a point where the original algorithms produce an almost identical path length distribution as the corrected ones.